# Raman Microspectroscopy for Real-Time Structure Indicator in Ultrafast Laser Writing

### v1

@@ VII Data availability
Supplementary Information
Supplementary Note 1: Pad optical micrograph and Raman map
@@
As desciribed in the manuscript, we employed laser-written graphitic pads and wires to emulate two common classes of conductive elements. Supplementary Figure 5a shows an optical micrograph of pad electrodes with a width of 50 m and length of 200 m, written at scan speeds of 100 and 200 m s-1, respectively. The 100 m s-1 pad exhibits a resistance of , whereas the 200 m s-1 pad shows , i.e., a difference of more than an order of magnitude. Under optical illumination, the laser-written tracks appear dark relative to the surrounding single-crystal diamond. This contrast arises because converting transparent, wide-band-gap sp3 diamond to disordered sp2 carbon increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [24, 25]. By contrast with the graphitic wires discussed in the main text, whose resistance difference is only weakly discernible in optical images, the two pads are optically indistinguishable. Supplementary Figures 5b and c present the corresponding photoluminescence (PL) and Raman (sp3 and sp2) maps. In both pads, the written regions appear bright in PL immediately after fabrication. This behavior differs from the wires, which appear dark in PL due to laser-induced non-radiative recombination pathways.
@@
The bright pad emission consists of locally activated NV0 and NV- centres in this Type Ib diamond (high nitrogen content), and other defects centres that producing a broad band that photo-bleaches under green (532 nm) excitation. This highlights that monitoring the full-window PL signal is not an ideal metric during fabrication, as relevant phase information is obscured by unrelated emissions and does not directly reflect device performance. In contrast, hyperspectral Raman mapping provides more specific structural information. By integrating the sp3 diamond window (1325-1340cm-1) [26] and the sp2 G-band window (1575-1610cm-1) [24], the sp3 map shows the laser-written regions as decreased peak intensity relative to pristine diamond, with stronger depletion for the lower scan speed. The sp2 map exhibits, while exhibiting the expected behavior (i.e., increased peak intensity in the written region), has lower overall contrast. Both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which arise from the writing process. Such features are more pronounced in the sp2 than in the sp3 maps. Details explanation will be discussed in Supplementary Note 2.
@@
Supplementary Figure 6 shows an optical micrograph of a pad acquired after PL mapping. Scanning a 20 mW, 532 nm beam around the electrode removes the initial femtosecond-laser-induced graphite debris [25], revealing a clean structure (dashed green square) compared to the post-writing state (dashed red square). Thus, maintaining the 532 nm probe during Raman acquisition could serve the dual role of (a) enabling spectroscopy monitoring, and (b) clearing graphite debris. With this, debris would not mask the written features during in situ monitoring, and no additional cleaning step is required.
Supplementary Note 2: Origin of contrast in hyperspectral maps
@@
As shown in Supplementary Figure 7, both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which are more pronounced in the sp2 map than in the sp3 map. The origin of these contrasts differs for the two cases. Since the pad is fabricated by stacking parallel wires with a 0.5 m pitch, the bright features in the sp3 maps correspond to regions between adjacent wires where less damage occurs, resulting in reduced graphitization. In contrast, the bright features in the sp2 maps arise from a different mechanism, consistent with the behavior described in Figure 1c of the main text. Here, the increased brightness is not related to enhanced graphitization but instead originates from broadband background emission.
@@
To confirm this, Supplementary Figure 7 presents a Raman map in the 590–600 nm spectral range (the same hyperspectral map of Figure 1c in the main text but different spectral window), excluding the sp3 and sp2 bands. Compared to the laser-written wire and the pure diamond background, the surrounding damaged shell appears bright, confirming that the observed contrast does not originate from either the sp3 or sp2 Raman signals. Spectra from three representative pixels—assigned to the graphite core, the damaged shell, and the pure diamond background—are extracted in Supplementary Figure 7a and shown in Supplementary Figure 7b–d. Each spectrum is self-normalized for comparison. The graphite region exhibits depleted sp3, a clear D band, and sp2 features, while the pure diamond spectrum is dominated by the sp3 line. In the damaged shell, no D or sp2 bands are observed apart from the sp3 peak. Instead, a broad emission is present, which contributes to the increased brightness observed in both the photoluminescence and hyperspectral maps. This emission is attributed to defect activation in the damaged shell surrounding the wire core [26], leading to an overall increase in count rates.
Supplementary Note 3: Pixel identification, integration, and normalization
@@
The hyperspectral Raman data was loaded as a four-dimensional array , where and represent the spatial coordinates, denotes the focal depth, and represents the wavelength. For all analyses, we selected the -plane closest to µm, effectively reducing the data to a three-dimensional array . To compensate for experimental variations including focusing, and excitation laser power fluctuations between different datasets, we implemented a sample-specific normalization procedure based on pristine diamond regions (shown as green pixels in Fig. 8).
@@
Supplementary Note 4: Quantitative PL analysis
@@
To evaluate whether broadband photoluminescence (PL) intensity can serve as an in situ process indicator—analogous to the sp3, sp2 signals and their ratio, we extracted the PL intensity from the written pixels. The results are presented in Supplementary Figure 9. This figure shows a correlation between the resistance and the normalized PL intensity of the written wire electrodes, which suggests that broadband PL intensity can be leveraged to assess performance of the written electrodes.
However, given the broad spectral collection window of these PL maps (550-850 nm), they can suffer from significant fluctuations due to factors unrelated to the phenomenon of interest, in this case, graphitization. For example, in Type Ib diamond, however, intrinsic defects generate a non-negligible background in PL maps; consequently, the brightest and darkest PL pixels do not directly correspond to “pure diamond” or “graphitized” regions.
Supplementary Figure 9, in this particular case, since the sp3 component dominates the broadband PL for the written region, the normalized PL follows the sp3 trend: PL increases with resistance, consistent with a larger fraction of diamond-like bonding. Nevertheless, the broadband PL also includes emission from defect clusters, leading to higher uncertainty (i.e. measurement variance) than for the sp3 metric.
@@
In summary, broadband PL can serve as a quantitative indicator during laser fabrication, but it is less specific and less accurate than the sp3 intensity. Accordingly, spectrally isolating the sp3 (and/or sp2) Raman windows—rather than relying on fully broadband PL—provides the most reliable and clean in situ process feedback.
Supplementary Note 5: sp3 distribution
@@
The speed-dependent probability-density histograms of the normalised sp3 Raman intensity, computed from pixels identified as graphitised in the hyperspectral maps, are shown for pad and wire electrodes in Supplementary Figure 10 and Supplementary Figure 11, respectively. For each scan speed, 1500 pixels are analysed for pads and 60 for wires. Intensities are referenced to pristine diamond, so higher normalised sp3 corresponds to weaker graphitization.
@@
In pad electrodes, increasing scan speed shifts the distributions toward higher values and simultaneously broadens them, consistent with reduced energy deposition per unit length and a less uniform sp3-to-sp2 transformation. By contrast, wire electrodes exhibit narrower distributions across all scan speeds because their geometry prevents overwriting. Nonetheless, they display the same overall trend as pads: increasing scan speed shifts the distributions to higher normalised sp3 values, reflecting a lower degree of graphitization in the laser-written structures.
Supplementary Note 6: sp3 Raman analysis
@@
To complement the two-probe transport data, we quantified how the sp3 Raman response varies with laser scan speed. For each speed, 60 pixels within the graphitized region of a laser-written wire were averaged and the sp3 peak was fitted with a Lorentzian. The wire geometry was chosen over the pad because it provides a cleaner, more localized transformation without overlapping features. Averaged spectra for scan speeds from 10 to 1270 m s-1 were analyzed; the resulting peak center and full width at half maximum (FWHM) are plotted in Fig. 12 as blue circles and dark-orange squares, respectively. A pristine-diamond reference, obtained by averaging 100 pixels from an unexposed area, is shown at 3000m s-1 for visual comparison only.
@@
Across the series, the sp3 mode lies slightly above 1332cm-1, consistent with tensile stress in the transformed graphite layer and the concomitant in-plane compressive stress imparted to the adjacent diamond[25], which produces a blue shift of the diamond Raman mode [31, 32]. The FWHM decreases systematically with increasing scan speed. Slower scans deposit more energy per unit length, driving a higher degree of graphitization; accordingly, the sp3 line departs from an ideal Lorentzian, broadening and sometimes becoming asymmetric. This behavior reflects increasing lattice disorder and strain distributions (inhomogeneous broadening), reduced phonon lifetimes (homogeneous broadening), and phonon-confinement effects; if the material becomes sufficiently conducting, Fano interference with an electronic continuum can further skew the lineshape [33, 47]. All averaged spectra and their Lorentzian fits are provided in Fig. 13.
Supplementary Note 7: Electrical performance for both pad and wire
@@
Supplementary Note 8: Hyperspectral unmixing
We performed hyperspectral unmixing with a linear mixing model using VCA for endmember extraction and FCLS for abundance estimation. As shown in Supplementary Figure 16, the unmixing results consistently show progressive loss of the sp3-like component with decreasing scan speed, while the sp2-like component increases and eventually saturates. This reinforces the main conclusion that sp3 depletion provides a reliable, monotonic indicator of graphitization. Hyperspectral unmixing is a powerful technique that leverages the full spectrum and does not require pre-defined labels. However, it is sensitive to the presence of noise or species not considered in the analysis. As shown in Supplementary Figure 16a, the presence of the broadband emission damage shell when fabricating at intermediate speeds is mislabeled as being a mixture of diamond and graphite phases.
Supplementary Figure 17 presents the endmembers extracted using different unmixing techniques and spectral regions of interest, exhibiting the variance that these endmembers can have. This behavior is expected: VCA and N-FINDR are geometric algorithms that approximate the data as lying within a convex simplex. They perform best when pixels lie close to the true extremes, but in practice the confocal voxel contains mixtures, variable SNR, NV0 fluorescence, and stress-induced peak shifts. The algorithms therefore identify different “extreme” representatives of the sp3-like and sp2-like responses, leading to small method-dependent variations that do not alter the physical interpretation. Indeed, the endmembers preserve the spectral features of their associated spectra, that is, the sp3 peak and NV0 band for the endmember associated to diamond, and the G band and, to lesser extent, the D band for the one associated to graphite.
Notably, for the ROI including the sp3 peak, the large dynamic range of this feature amplifies small fluctuations which can distort the convex geometry and bias endmember selection. Excluding this line (ROI 1345–1800 cm-1) reduces leverage and produces noisier but qualitatively similar maps.
It is important to distinguish geometric unmixing from PCA. VCA and N-FINDR seek spectra at the boundaries of the data cloud that can be interpreted as approximate endmembers, whereas PCA decomposes variance into orthogonal components that need not resemble real spectra and can take negative values. PCA is valuable for denoising and diagnostics, but its components cannot be interpreted as abundances.

### v2

@@ 7 Data availability
Supporting Information
Supporting Note 1: Pad optical micrograph and Raman map
@@
As described in the manuscript, we employed laser-written graphitic pads and wires to emulate two common classes of conductive elements. Supporting Figure 1a shows an optical micrograph of pad electrodes with a width of 50 m and length of 200 m, written at scan speeds of 100 and 200 m s-1, respectively. The 100 m s-1 pad exhibits a resistance of , whereas the 200 m s-1 pad shows , i.e., a difference of more than an order of magnitude. Under optical illumination, the laser-written tracks appear dark relative to the surrounding single-crystal diamond. This contrast arises because converting transparent, wide-band-gap sp3 diamond to disordered sp2 carbon increases visible-wavelength absorption and slightly raises the refractive index, thereby reducing transmittance through the modified region [prelas1997handbook, Wang2000]. Supporting Figures 1b and c present the corresponding photoluminescence (PL) and Raman (sp3 and sp2) maps. In both pads, the written regions appear bright in PL immediately after fabrication. This behavior differs from the wires, which appear dark in PL due to laser-induced non-radiative recombination pathways.
@@
The bright pad emission consists of locally activated NV0 and NV- centers in this Type Ib diamond (high nitrogen content), and other defects centers that producing a broad band that photo-bleaches under green (532 nm) excitation. This highlights that monitoring the full-window PL signal is not an ideal metric during fabrication, as relevant phase information is obscured by unrelated emissions and does not directly reflect device performance. In contrast, hyperspectral Raman mapping provides more specific structural information. By integrating the sp3 diamond window (1325-1340cm-1) [Kononenko2008] and the sp2 G-band window (1575-1610cm-1) [prelas1997handbook], the sp3 map shows the laser-written regions as decreased peak intensity relative to pristine diamond, with stronger depletion for the lower scan speed. The sp2 map exhibits the expected behavior (i.e., increased peak intensity in the written region) but has lower overall contrast. Both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which arise from the writing process. Such features are more pronounced in the sp2 than in the sp3 maps. Details explanation will be discussed in Supporting Note 2.
@@
Supporting Figure 2 shows an optical micrograph of a pad acquired after PL mapping. Scanning a 20 mW, 532 nm beam around the electrode removes the initial femtosecond-laser-induced graphite debris [Wang2000], revealing a clean structure (dashed green square) compared to the post-writing state (dashed red square). Thus, maintaining the 532 nm probe during Raman acquisition could serve the dual role of (a) enabling spectroscopy monitoring, and (b) clearing graphite debris.
Supporting Note 2: Origin of contrast in hyperspectral maps
@@
As shown in Supporting Figure 3, both the sp3 and sp2 Raman maps exhibit horizontal line-like features, which are more pronounced in the sp2 map than in the sp3 map. The origin of these contrasts differs for the two cases. Since the pad is fabricated by stacking parallel wires with a 0.5 m pitch, the bright features in the sp3 maps correspond to regions between adjacent wires where less damage occurs, resulting in reduced graphitization. In contrast, the bright features in the sp2 maps arise from a different mechanism, consistent with the behavior described in Figure 1c of the main text. Here, the increased brightness is not related to enhanced graphitization but instead originates from broadband background emission.
@@
To confirm this, Supporting Figure 3 presents a Raman map in the 590–600 nm spectral range (the same hyperspectral map of Figure 1c in the main text but different spectral window), excluding the sp3 and sp2 bands. Compared to the laser-written wire and the pure diamond background, the surrounding damaged shell appears bright, confirming that the observed contrast does not originate from either the sp3 or sp2 Raman signals. Spectra from three representative pixels—assigned to the graphite core, the damaged shell, and the pure diamond background—are extracted in Supporting Figure 3a and shown in Supporting Figure 3b–d. Each spectrum is normalized for comparison. The graphite region exhibits depleted sp3, a clear D band, and sp2 features, while the pure diamond spectrum is dominated by the sp3 line. In the damaged shell, no D or sp2 bands are observed apart from the sp3 peak. Instead, a broad emission is present, which contributes to the increased brightness observed in both the photoluminescence and hyperspectral maps. This emission is attributed to defect activation in the damaged shell surrounding the wire core [Kononenko2008].
Supporting Note 3: Pixel identification, integration, and normalization
@@
The hyperspectral Raman data was loaded as a three-dimensional array , where and represent the spatial coordinates, and represents the wavelength. To compensate for experimental variations including focusing, and excitation laser power fluctuations between different datasets, we implemented a sample-specific normalization procedure based on pristine diamond regions (shown as green pixels in Fig. 4).
@@
Supporting Note 4: Quantitative PL analysis
@@
To evaluate whether broadband photoluminescence (PL) intensity can serve as an in situ process indicator, we extracted the PL intensity from the written pixels. The results are presented in Supporting Figure 5. This figure shows a correlation between the resistance and the normalized PL intensity of the written wire electrodes, which suggests that broadband PL intensity can be leveraged to assess performance of the written electrodes.
Supporting Figure 5 also compares the PL with the sp3 intensity metric proposed in the main manuscript. This comparison showcases how the normalized PL follows the sp3 trend: PL increases with resistance, consistent with a larger fraction of diamond-like bonding. Yet, despite having similar correlation with the resistance of the electrodes, the PL normalized intensity presents significantly larger variance in the measurements. This is due to the broad spectral collection window of these PL maps (550-850 nm), which can include emission from phenomena not directly related to graphitization (i.e., in Type Ib diamond intrinsic defects generate a non-negligible background in PL maps).
@@
Thus, even if broadband PL can serve as a quantitative indicator during laser fabrication, in practice it is less specific and less accurate than the sp3 intensity, which should be favored for in situ process monitoring.
Supporting Note 5: sp3 distribution
@@
The speed-dependent probability-density histograms of the normalized sp3 Raman intensity, computed from pixels identified as graphitized in the hyperspectral maps, are shown for pad and wire electrodes in Supporting Figure 6 and Supporting Figure 7, respectively. For each scan speed, 1500 pixels are analyzed for pads and 60 for wires. Intensities are referenced to pristine diamond, so higher normalized sp3 corresponds to weaker graphitization.
@@
In pad electrodes, increasing scan speed shifts the distributions toward higher values and simultaneously broadens them, consistent with reduced energy deposition per unit length and a less uniform sp3-to-sp2 transformation. By contrast, wire electrodes exhibit narrower distributions across all scan speeds because their geometry prevents overwriting. Nonetheless, they display the same overall trend as pads: increasing scan speed shifts the distributions to higher normalized sp3 values, reflecting a lower degree of graphitization in the laser-written structures.
Supporting Note 6: sp3 Raman analysis
@@
To complement the two-probe transport data, we quantified how the sp3 Raman response varies with laser scan speed. For each speed, 60 pixels within the graphitized region of a laser-written wire were averaged and the sp3 peak was fitted with a Lorentzian. Averaged spectra for scan speeds from 10 to 1270 m s-1 were analyzed; the resulting peak center and full width at half maximum (FWHM) are plotted in Fig. 8 as blue circles and dark-orange squares, respectively. A pristine-diamond reference, obtained by averaging 100 pixels from an unexposed area, is shown at 3000m s-1 for visual comparison only.
@@
Across the series, the sp3 mode lies slightly above 1332cm-1, consistent with tensile stress in the transformed graphite layer and the concomitant in-plane compressive stress imparted to the adjacent diamond[Wang2000], which produces a blue shift of the diamond Raman mode [Grimsditch1975, Zaitsev2001]. The FWHM decreases systematically with increasing scan speed. Slower scans deposit more energy per unit length, driving a higher degree of graphitization; accordingly, the sp3 line departs from an ideal Lorentzian, broadening and sometimes becoming asymmetric. This behavior reflects increasing lattice disorder and strain distributions (inhomogeneous broadening), reduced phonon lifetimes (homogeneous broadening), and phonon-confinement effects; if the material becomes sufficiently conducting, Fano interference with an electronic continuum can further skew the lineshape [Orwa2000, Orwa2009]. All averaged spectra and their Lorentzian fits are provided in Fig. 9.
Supporting Note 7: Electrical performance for both pad and wire
@@
Supporting Note 8: Hyperspectral unmixing
We performed hyperspectral unmixing with a linear mixing model using vertex component analysis (VCA) [nascimento2005vertex] for endmember extraction and fully-constrained least squares (FCLS) for abundance estimation. As shown in Supporting Figure 12, the unmixing results consistently show progressive loss of the diamond-associated component with decreasing scan speed, while the carbon-associated component increases and eventually saturates. This reinforces the main conclusion that sp3 depletion provides a reliable, monotonic indicator of graphitization. Hyperspectral unmixing is a powerful technique that leverages the full spectrum and does not require pre-defined labels. However, it is sensitive to the presence of noise or species not considered in the analysis. As shown in Supporting Figure 12a, the presence of the broadband emission damage shell when fabricating at intermediate speeds is mislabeled as being a mixture of diamond and graphite phases, since in the analysis only two endmembers are considered.
Supporting Figure 13 presents the endmembers extracted using different unmixing techniques and spectral regions of interest (ROI), exhibiting the variance that these endmembers can have. This behavior is expected: VCA and N-FINDR [winter1999n] are geometric algorithms that approximate the data as lying within a convex simplex. They perform best when pixels lie close to the true extremes, but in practice the confocal voxel contains mixtures, variable SNR, NV0 fluorescence, and potentially stress-induced peak shifts. The algorithms therefore identify different “extreme” representatives of the diamond-like and graphite-like responses, leading to small method-dependent variations that do not alter the physical interpretation. Indeed, the endmembers preserve the spectral features of their associated spectra, that is, the sp3 peak and NV0 band for the endmember associated to diamond, and the G band and, to lesser extent, the D band for the one associated to graphite.
Notably, for the spectral ROI including the sp3 peak, the large dynamic range of this feature amplifies small fluctuations which can distort the convex geometry and bias endmember selection. Excluding this line (spectral ROI: 1345–1800 cm-1) reduces leverage and produces noisier but qualitatively similar maps.
Supporting Figure 13 also provides the first two principal components. It is important to distinguish geometric unmixing from principal component analysis (PCA) [abdi2010principal]. VCA and N-FINDR seek spectra at the boundaries of the data cloud that can be interpreted as approximate endmembers, whereas PCA decomposes variance into orthogonal components that need not resemble real spectra and can take negative values. PCA is valuable for denoising and diagnostics, but its components cannot be interpreted as abundances.